Solve for x. Round to the nearest ten-thousandth
log (538x) = 3.631

Answer by Alan3354(21550)

log (538x) = 3.631
---------
10^(log(538x)) = 10^3.631
538x = 10^3.631 = 4275.62886
x = 4275.../538
x = 7.9473